A refrigerator has a coefficient of performance of 1.15, and it extracts 7.95 J of heat from the
cold reservoir during each cycle. (a) How much work is done on the gas in each cycle? (b)
How much heat is exhausted into the hot reservoir in each cycle?

I've seen a similar question, but I still don't understand how to solve. Can I get help please, step by step would be kind.
Answer #1

a)

from definition of COP

1.15 = Q/W

W = 7.95 / 1.15 = 6.91 J

=====

b)

now,

Q' = Q + h

Q'= 7.95 + 6.91 = 14.86 J